OpenStudy (anonymous):

10 ?? ___ = _____ 15 27 WILL FAN AND MEDAL

HanAkoSolo (jamierox4ev3r):

this is a ratio. All you have to do to find the variable in question is set it as "x" and cross multiply. \(\Large\frac{10}{15}\) = \(\Large\frac{x}{27}\) Then, cross multiply 15x=270 Make sense so far?

OpenStudy (anonymous):

Kind of. (I'm out of practice with math...)

OpenStudy (anonymous):

Wait, no I see now. I mis read

HanAkoSolo (jamierox4ev3r):

That's fine. So once you have 15x=270, all you have to do solve for x. and you do that by dividing both sides by 15. Do you think you could solve this?

OpenStudy (anonymous):

so 18/27?

HanAkoSolo (jamierox4ev3r):

That's exactly right. The "??" would be equal to 18. so 10/15 is equal to 18/27. I hope this makes sense to you :)

OpenStudy (anonymous):

It makes perfect sense thank you so much (:

HanAkoSolo (jamierox4ev3r):

Any time :)

hero (hero):

Don't forget to simplify first if possible. Working with the smallest numbers possible helps make the multiplications easier. 10/15 simplifies to 2/3 so you end up with 2/3 = x/27 Which cross multiplies to 54 = 3x

OpenStudy (anonymous):

Thanks (: